if John is 3 times older than Mary and there is a 24 year difference in their ages, how old are they?
Found 2 solutions by htmentor, jorel1380:
Answer by htmentor(1343)

Let J = John's age
Let M = Mary's age
Given:
J = 3M
J - M = 24
Substitute the value for J in the 2nd equation:
3M - M = 24
Solving for M gives M = 12
Therefore J = 3*12 = 36
Ans: John = 36 yrs old, Mary = 12 yrs old

Answer by jorel1380(3719)

Let J equal John, M for Mary;
J=3M
J-M=24
Substituting for J we get;
3M-M=24
2M=24
M=12
Mary is 12.
Since J=3m, John is 36.